y^{'}+6y=4e^{-5x}


1
Expert's answer
2021-05-07T09:41:51-0400

y^{'}+6y=4e^{-5x}

The problem is rewritten as

"\frac{dy}{dx} + 6y = 4e^{-5x}"

This is a first order linear differential equation

Integrating factor = "e^{\smallint6dx} = e^{6x}"

Multiplying both sides by the integrating factor we get

"e^{6x}[\frac{dy}{dx} + 6y] = 4e^{-5x}e^{6x}"

=> "e^{6x}\frac{dy}{dx} + 6e^{6x}y = 4e^{x}"

=> "e^{6x}\frac{dy}{dx} + \frac{de^{6x}}{dx}y = 4e^{x}"

=> "\frac{d(ye^{6x})}{dx} = 4e^{x}"

=> "d(ye^{6x})= 4e^{x}dx"

Integrating both sides

"\int{d(ye^{6x})}=\int{ 4e^{x}dx}"

=> "ye^{6x} = 4e^{x}+C" where C is integration constant